Steam enters with negligible velocity a well-insulated turbine operating at 

steady state at 4 MPa, 320°C. The steam expands to an exit pressure of 0.07 

MPa, a quality (dryness fraction) of 0.95 and a velocity of 90 m/s. The diameter 

of the exit is 0.6 m. Neglecting potential energy effects, find the power 

developed by the turbine in kW



1
Expert's answer
2021-07-08T13:17:24-0400

Gives


"P_1=4MPa,T_1=320\u00b0c,P_2=0.07Mpa,\nv=90m\/sec"

r=0.3m

"T_2=27\u00b0c"

Quality=0.95

"\dot{m}=Av"


"A=\pi r^2=3.14\times0.3^2=0.2826m^2"

"\dot{m}=0.2826\times90=25.434kg/sec"

Power

"P=\frac{dQ}{dt}"

We know that

"\dot{Q}=\dot{m}c_p∆T"


"\frac{dQ}{dt}=25.43\times4.18(320-27)=31145.13J/sec"

Now

"P=31.145KJ\/sec"